# The Meta-Representation Hypothesis

### v3

@@ Appendix A Proofs
Lemma A.1 (Performance difference).
@@
This result can be directly derived from Kakade & Langford (2002). ∎
Lemma A.2.
@@
$\begin{split}\left|\mathbb{E}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]\right|&=\left|\mathbb{E}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]-\mathbb{E}_{a\sim\mu(\cdot|s)}\left[A^{\mu}(s,a)\right]\right|\\ &\leq\left\|\tilde{\mu}(a|s)-\mu(a|s)\right\|_{1}\cdot\left\|A^{\mu}(s,a)\right\|_{\infty}\\ &=2D_{\mathrm{TV}}(\tilde{\mu}\|\mu)[s]\cdot\max_{a}\left|A^{\mu}(s,a)\right|.\\ \end{split}$ (27)
This is a widely used trick (Schulman, 2015; Zhuang et al., 2023; Gan et al., 2024). ∎
@@
Next, according to Lemma A.2, we have
$\left\|d^{\tilde{\mu}}-d^{\mu}\right\|_{1}\cdot\left\|\mathop{\mathbb{E}}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]\right\|_{\infty}=\epsilon\left\|d^{\tilde{\mu}}-d^{\mu}\right\|_{1}\leq\frac{2\epsilon\gamma}{1-\gamma}\mathop{\mathbb{E}}_{s\sim d^{\mu}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}\|\mu)[s]\right],$ (31)
@@
Let’s start with the first-order approximation of the training performance (Schulman, 2015), denote it as
@@
Then, we are trying to bound the difference between $\zeta(\tilde{\pi})$ and $L_{\pi}(\tilde{\pi})$, according to Lemma A.1, that is,
@@
$\begin{split}&\left|\int_{\mathcal{F}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}p_{\mathrm{train}}(f)\cdot g(f)\mathrm{d}f\right|\\ =&\left|\int_{\mathcal{F}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot g(f)\mathrm{d}f\right|\\ =&\left|\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot g(f)\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f\right|\\ =&\left|\int_{\mathcal{F}_{\mathrm{train}}}\frac{Z-1}{Z}p(f)\cdot g(f)\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f\right|,\\ \end{split}$ (37)
@@
Using Corollary A.4, $\Psi$ can be bounded by
@@
where $\epsilon=\max_{s}\left|\mathbb{E}_{a\sim\tilde{\mu}_{f}(\cdot|s)}\left[A^{\mu_{f}}(s,a)\right]\right|$, denote $\epsilon_{\mathrm{train}}=\max_{f\in\mathcal{F}_{\mathrm{train}}}\left\{\max_{s}\left|\mathbb{E}_{a\sim\tilde{\mu}_{f}(\cdot|s)}\left[A^{\mu_{f}}(s,a)\right]\right|\right\}$, we obtain
@@
According to Lemma A.3, we can bound $g(f)$, which can be expressed as
@@
$\begin{split}\Phi&\leq\frac{1-Z}{Z}\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot\left|g(f)\right|\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot\left|g(f)\right|\mathrm{d}f\\ &\leq\frac{2\delta_{\mathrm{train}}(1-Z)}{Z}\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f+2\delta_{\mathrm{eval}}\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ &=2\delta_{\mathrm{train}}(1-Z)\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f+2\delta_{\mathrm{eval}}(1-Z)\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{1-Z}\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ &=2\delta_{\mathrm{train}}(1-Z)\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]+2\delta_{\mathrm{eval}}(1-Z)\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{eval}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right],\\ \end{split}$ (50)
@@
According to Lemma A.2, we have
@@
Similar to the proof of Theorem 3.4, using Lemma A.2 again, we have
@@
Appendix B A More Detailed Explanation of Our Hypothesis
In Section 4, we claimed that “The DML loss encourages them to make consistent decisions on the same observations, meaning that any irrelevant features learned by policy A are likely to result in suboptimal performance for policy B, and vice versa.” Here, we aim to provide a more detailed explanation to help readers better understand this point.
Let’s consider a simple environment where the agent is in a rectangular space and attempts to pick up coins to earn rewards (see Figure 6). The agent’s observations are the current pixels.
Figure 6: This is a simple rectangular environment where the gray agent’s goal is to pick up circular coins.
It is clear that the agent’s true objective is to pick up the coins, and the background color is a spurious feature. However, upon observing the training data for policy A, we can see that in the red background, the coins are always on the right side of the agent, while in the cyan background, the coins are always on the left side. As a result, when training policy A using reinforcement learning algorithms, it is likely to exhibit overfitting behavior, such as moving to the right in a red background and to the left in a cyan background.
However, the overfitting of policy A to the background color will fail in the training data of policy B, because in policy B’s training data, regardless of whether the background color is red or cyan, the coin can appear either on the left or right side of the agent. Therefore, through DML, policy A is regularized by the behavior of policy B during the training process, effectively preventing policy A from overfitting to the background color. In other words, any irrelevant features learned by policy A could lead to suboptimal performance of policy B, and vice versa. Thus, we hypothesize that this process will force both policy A and policy B to learn the true underlying semantics, ultimately converging to meta-representations.
Appendix C More Empirical Results
Appendix D More Implementation Details
D.1 Proximal Policy Optimization
In our experiments, we employ Proximal Policy Optimization (PPO) as our baseline algorithm. Specifically, given the policy network $\pi_{\theta}$, the value network $V_{\phi}$, and any observation-action pair $(o_{t},a_{t})$, the value loss is
$\mathcal{L}_{v}=\frac{1}{2}[V_{\phi}(o_{t})-\hat{R}_{t}]^{2},$ (78)
where $\hat{R}_{t}$ is the estimated discounted return at step $t$ using the Generalized Advantage Estimation (GAE) (Schulman et al., 2015) technique. And the policy loss is
$\mathcal{L}_{p}=-\min\left[\frac{\pi_{\theta}(a_{t}|o_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})}\cdot\hat{A}(o_{t},a_{t}),\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}|o_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})},1-\epsilon,1+\epsilon\right)\cdot\hat{A}(o_{t},a_{t})\right],\enspace\mathcal{L}_{e}=\mathcal{H}(\pi_{\theta}(\cdot|o_{t})),$ (79)
where $\mathcal{H}(\cdot)$ represents the entropy of the output action distribution. The pseudo-code for PPO is provided in Algorithm 3.
Algorithm 3 Proximal Policy Optimization (PPO)
1:  Initialize: Policy and value nets $\pi_{\theta}$ and $V_{\phi}$, clipping parameter $\epsilon$, value loss coefficient $c_{1}$, policy entropy coefficient $c_{2}$
2:  Output: Optimal policy network $\pi_{\theta^{*}}$
3:  while not converged do
4:     # Data collection
5:     Collect data $\mathcal{D}=\left\{(o_{t},a_{t},r_{t})\right\}_{t=1}^{N}$ using the current policy network $\pi_{\theta}$
6:     # The networks before updating
7:     $\pi_{\theta_{\mathrm{old}}}\leftarrow\pi_{\theta},\enspace V_{\phi_{\mathrm{old}}}\leftarrow V_{\phi}$
8:     # Estimate the advantage $\hat{A}(o_{t},a_{t})$ based on $V_{\phi_{\mathrm{old}}}$
9:     Use GAE (Schulman et al., 2015) technique to estimate the advantage $\hat{A}(o_{t},a_{t})$
10:     # Estimate the return $\hat{R}_{t}$
11:     $\hat{R}_{t}\leftarrow V_{\phi_{\mathrm{old}}}(o_{t})+\hat{A}(o_{t},a_{t})$
12:     for each training epoch do
13:        # Compute policy loss $\mathcal{L}_{p}$
14:        $\mathcal{L}_{p}\leftarrow-\frac{1}{N}\sum_{t=1}^{N}\min\left[\frac{\pi_{\theta}(a_{t}|o_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})}\cdot\hat{A}(o_{t},a_{t}),\mathrm{clip}\left(\frac{\pi_{\theta}(a_{t}|o_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}|o_{t})},1-\epsilon,1+\epsilon\right)\cdot\hat{A}(o_{t},a_{t})\right]$
15:        # Compute policy entropy $\mathcal{L}_{e}$ and value loss $\mathcal{L}_{v}$
16:        $\mathcal{L}_{e}\leftarrow\frac{1}{N}\sum_{t=1}^{N}\mathcal{H}(\pi_{\theta}(\cdot|o_{t})),\enspace\mathcal{L}_{v}\leftarrow\frac{1}{2N}\sum_{t=1}^{N}[V_{\phi}(o_{t})-\hat{R}_{t}]^{2}$
17:        # Compute total loss $\mathcal{L}$
18:        $\mathcal{L}\leftarrow\mathcal{L}_{p}+c_{1}\mathcal{L}_{v}-c_{2}\mathcal{L}_{e}$
19:        # Update parameters $\theta$ and $\phi$ through backpropagation, $\lambda_{\theta}$ and $\lambda_{\phi}$ is the step sizes
20:        $\theta\leftarrow\theta-\lambda_{\theta}\nabla_{\theta}\mathcal{L},\enspace\phi\leftarrow\phi-\lambda_{\phi}\nabla_{\phi}\mathcal{L}$
21:     end for
22:  end while
D.2 PPO with DML
Our approach introduces an additional KL divergence loss to encourage mutual learning between the two agents, which is
$\mathcal{L}_{\mathrm{DML}}=\mathcal{L}_{p}+c_{1}\mathcal{L}_{v}-c_{2}\mathcal{L}_{e}+\alpha\mathcal{L}_{\mathrm{KL}},$ (80)
where $\mathcal{L}_{p}+c_{1}\mathcal{L}_{v}-c_{2}\mathcal{L}_{e}$ is the reinforcement learning loss, and
$\mathcal{L}_{\mathrm{KL}}=D_{\mathrm{KL}}(\pi_{\hat{\theta}}\|\pi_{\theta})$ (81)
is the KL divergence between the current policy and the other agent’s policy, $\alpha$ is the weight, and $\pi_{\hat{\theta}}$ denotes the other agent’s policy. Thus, this additional KL loss encourages the two agents to make consistent decisions for the same observations.
D.3 Hyperparameter Settings
Table 2 shows the detailed hyperparameter settings in our code, with the main hyperparameters consistent with the hard-level settings in Cobbe et al. (2020), except that we trained for 50M steps instead of 200M. We trained the policy on the initial 500 levels and tested its generalization performance across the entire level distribution.
Table 2: Detailed hyperparameters in Procgen.
[…]
Number of workers 64 64
Horizon 256 256
Learning rate 0.0005 0.0005
Learning rate decay No No
Optimizer Adam Adam
Total interaction steps 50M 50M
Update epochs 3 3
Mini-batches 8 8
Batch size 16384 16384
Mini-batch size 2048 2048
Discount factor $\gamma$ 0.999 0.999
GAE parameter $\lambda$ 0.95 0.95
Value loss coefficient $c_{1}$ 0.5 0.5
Entropy loss coefficient $c_{2}$ 0.01 0.01
Clipping parameter $\epsilon$ 0.2 0.2
KL divergence weight $\alpha$ 1.0

### v4

@@ Appendix G Proofs
Lemma G.1 (Performance difference).
@@
This result can be directly derived from Kakade and Langford [20]. ∎
Lemma G.2.
@@
$\begin{split}\left|\mathbb{E}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]\right|&=\left|\mathbb{E}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]-\mathbb{E}_{a\sim\mu(\cdot|s)}\left[A^{\mu}(s,a)\right]\right|\\ &\leq\left\|\tilde{\mu}-\mu\right\|_{1}\cdot\left\|A^{\mu}(s,a)\right\|_{\infty}\\ &=2D_{\mathrm{TV}}(\tilde{\mu}\|\mu)[s]\cdot\max_{a}\left|A^{\mu}(s,a)\right|.\\ \end{split}$ (30)
This is a widely used trick [36, 55, 12]. ∎
@@
Next, according to Lemma G.2, we have
$\left\|d^{\tilde{\mu}}-d^{\mu}\right\|_{1}\cdot\left\|\mathop{\mathbb{E}}_{a\sim\tilde{\mu}(\cdot|s)}\left[A^{\mu}(s,a)\right]\right\|_{\infty}=\left\|d^{\tilde{\mu}}-d^{\mu}\right\|_{1}\cdot\epsilon\leq\frac{2\epsilon\gamma}{1-\gamma}\mathop{\mathbb{E}}_{s\sim d^{\mu}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}\|\mu)[s]\right],$ (34)
@@
Let’s start with the first-order approximation of the training performance [36], denote it as
@@
Then, we are trying to bound the difference between $\zeta(\tilde{\pi})$ and $L_{\pi}(\tilde{\pi})$, according to Lemma G.1, that is,
@@
$\begin{split}&\left|\int_{\mathcal{F}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}p_{\mathrm{train}}(f)\cdot g(f)\mathrm{d}f\right|=\left|\int_{\mathcal{F}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot g(f)\mathrm{d}f\right|\\ =&\left|\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f-\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot g(f)\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f\right|\\ =&\left|\int_{\mathcal{F}_{\mathrm{train}}}\frac{Z-1}{Z}p(f)\cdot g(f)\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot g(f)\mathrm{d}f\right|,\\ \end{split}$ (40)
@@
Using Corollary G.4, $\Psi$ can be bounded by
@@
where $\epsilon=\max_{s}\left|\mathbb{E}_{a\sim\tilde{\mu}_{f}(\cdot|s)}\left[A^{\mu_{f}}(s,a)\right]\right|$, denote $\epsilon_{\mathrm{train}}=\max_{f\in\mathcal{F}_{\mathrm{train}}}\left\{\epsilon\right\}$, we obtain
@@
According to Lemma G.3, we can bound $g(f)$, which can be expressed as
@@
$\begin{split}\Phi\leq&\frac{1-Z}{Z}\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot\left|g(f)\right|\mathrm{d}f+\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot\left|g(f)\right|\mathrm{d}f\\ \leq&\frac{2\delta_{\mathrm{train}}(1-Z)}{Z}\int_{\mathcal{F}_{\mathrm{train}}}p(f)\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ &+2\delta_{\mathrm{eval}}\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}p(f)\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ =&2\delta_{\mathrm{train}}(1-Z)\int_{\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{Z}\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ &+2\delta_{\mathrm{eval}}(1-Z)\int_{\mathcal{F}-\mathcal{F}_{\mathrm{train}}}\frac{p(f)}{1-Z}\cdot\mathop{\mathbb{E}}_{s\sim d^{\tilde{\mu}_{f}}(\cdot)}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]\mathrm{d}f\\ =&2\delta_{\mathrm{train}}(1-Z)\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{train}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right]+2\delta_{\mathrm{eval}}(1-Z)\mathop{\mathbb{E}}_{\begin{subarray}{c}f\sim p_{\mathrm{eval}}(\cdot)\\ s\sim d^{\tilde{\mu}_{f}}(\cdot)\end{subarray}}\left[D_{\mathrm{TV}}(\tilde{\mu}_{f}\|\mu_{f})[s]\right],\\ \end{split}$ (53)
@@
According to Lemma G.2, we have
@@
Similar to the proof of Theorem 3.4, using Lemma G.2 again, we have